Page tree
Skip to end of metadata
Go to start of metadata


Attendance: David Williams, Tim Grundy, Scott Wallace, Radford Perry.

Apologies sent: Udo Schühle.


CMS ops

Currently executing on board is STP-241. This STP is part of period known as LTP10, which runs from until .

Anomalies (if any)

On at around 12:30 UT, during scheduled maintenance of the Solid State Mass Memory, there was an unexpected failure when commanding Supervisor-B from INIT to OPERA, after patching Supervisor-A (in SERVICE mode), and switching SV-A OFF.

  • SPICE was consequently switched OFF.
  • The FCT was given the go ahead to cold-boot the whole SSMM unit, as no significant science data were stored in the mass memory.
  • The MTL was disabled so no instrument commands would be sent to the payload even after the end of the engineering window at 17:00Z.

SPICE was restarted around at 13:00 UT the following day , and brought back to the original timeline by 17:00 UT.

TG: all date from 16 and 17 Jan were already down, including last set of CMS data. Lost science data from 18th Jan a.m. 

(SPICE weekly calibration data were lost, though.)

CMS data and interpretation 

See IAS CMS data tool
 

All plots below are clickable to zoom in, and can be directly downloaded from this page.


Operations times shown for CMS since the last meeting.

LTP09:

  1. STP-234
    • 2022-12-05T00:05:00Z – 2022-12-05T03:05:00Z
      • WOL
    • 2022-12-09T12:41:50Z – 2022-12-09T22:41:50Z
      • TCM (Type-2) relatively large: v = 5 m/s. 1803 – 1820 UT

      • Thermal control was used on CMS-2 (thanks to Tim for the idea) set at 5ºC (measured at 5.2ºC).

      • Seem to pick up 1.5 Hz in frequency change on CMS-2
      • Similar temp before and after, although slightly higher afterwards. 2.5695 → 2.5701 about +0.6 Hz residual. 
  2. STP-235
    • 2022-12-12T00:05:00Z – 2022-12-12T03:05:00Z
      • WOL
    • 2022-12-17T07:10:00Z – 2022-12-17T10:10:00Z
      • SA_ROT (30º → 0º)
  3. STP-236
    • 2022-12-19T00:00:00Z – 2022-12-20T00:00:00Z
      • WOL, TCM  and MAINT (Any)
    • 2022-12-20T10:00:00Z – 2022-12-20T13:00:00Z
      • SA Characterisation 11:00 – 11:28
        • Would be good to see if this lets us fill in any missing angle-angle effects on the SA.
        • ~5-minute offset between highest SA inclination angle and the minimum temperature.
        • If we can pull out the time constant of the thermal changes from the data we already have from regular SA_ROT events, this could be a really useful input dataset.
  4. STP-237
    • 2022-12-26T00:05:00Z – 2022-12-26T03:05:00Z
      • WOL

LTP-10

  1. STP-238
    • 2023-01-02T00:05:00Z – 2023-01-02T03:05:00Z
      • WOL
  2. STP-239
    • 2023-01-09T00:05:00Z – 2023-01-09T03:05:00Z
      • WOL
    • 2023-01-12T22:00:00Z – 2023-01-13T01:00:00Z
      • SA Relubrication
  3. STP-240
    • 2023-01-16T00:00:00Z – 2023-01-17T00:00:00Z
      • WOL
      • TCM (type-2)
      • MAINT-ANY (beyond 0.95 AU, so off-pointing from sun centre is possible).
    • Maybe a small jump in frequency there, but needs to be visible in MISO, so there was a thruster-induced blip.

Future operations

LTP10

  1. STP-241
    • 2023-01-23T00:05:00Z – 2023-01-23T03:05:00Z
      • WOL
  2. STP-242
    • 2023-01-30T00:05:00Z – 2023-01-30T03:05:00Z
      • WOL
    • 2023-02-04T13:00:00Z – 2023-02-04T16:00:00Z
      • SA_ROT (30º → 56º)
  3. STP-243
    • 2023-02-06T00:05:00Z – 2023-02-06T03:05:00Z
      • WOL
  4. STP-244
    • 2023-02-13T00:00:00Z – 2023-02-14T00:00:00Z
      • WOL
      • TCM (Type-2)
      • MAINT-ANY (inside 0.95 AU)
  5. STP-245
    • 2023-02-20T00:05:00Z – 2023-02-20T03:05:00Z
      • WOL


Being planned

LTP11

This plan is soon to be released. SPICE will have the option to check timings for EMC-related conflicts before it's confirmed.

To be planned

LTP12

Will need to plot the next TGA for early July.



Long-term trends


Recommendations 

  • CMS operations to be performed weekly, capturing interesting events where possible.
  • CMS Operator should switch on CMS ≥30 minutes earlier than the time at which the observations would start, particularly if the observations to catch (for example) a WOL would need to start at midnight. CMS operations in SOOP Kitchen are timed such that they start 30 minutes after assumed switch-on.
  • Note that From LTP04 onwards (i.e., starting from STP-154), SOC has not been indicating CMS operations in the EFECS with an RSW_EXT_SPICE. They will still be indicated with CMS “observations” in the LTP plan however.
  • For MAINT-ANY windows, since FCT can do ANYTHING in that window, it's good practice to observe right through (and for any adjoining WOLs + TCMs).


Tasks for the near- to mid-term future

YGT proposal wasn't possible this time around, but we can look out for the next opportunities. Could involve TEC, too. Idea was to have them work out the cumulative dose from SWA, EPD, etc.  CReMA had some radiation requirements, orbital analysis, shielding requirements.



AOBs


Next meeting

, proposing to return to our old schedule of the last Thursday of each month at 17:00 CET (Subject to confirmation on Monday )





Action Items

  • No labels